### v1

# Seal: Symbolic Execution with Separation Logic

Abstract
Seal is a static analyser for the verification of programs that manipulate unbounded linked data structures. It is based on separation logic to represent abstract memory states and, unlike other separation-logic-based approaches, it employs a general-purpose separation logic solver Astral for satisfiability and entailment checking, which itself is based on translation to SMT. This design results in a modular architecture intended to be easier to extend and to combine with reasoning in other theories. Although still a prototype, Seal achieved competitive results in the LinkedLists base category and was one of only four analysers capable of verifying programs with unbounded lists. We believe that the tool’s extensibility, combined with further development, can lead to significant improvements in future competitions.
1 Verification Approach
Seal is a shape analyser based on forward abstract interpretation over formulae of separation logic (SL) [19], inspired by the approach of [12]. Unlike the vast majority of SL-based analysers, Seal relies on a general-purpose SL solver Astral [11] to check satisfiability and entailment of SL formulae (instead of using some dedicated heuristic approach for this purpose). As a result, analysis and logical reasoning in SL are cleanly decoupled. The Astral solver was originally developed by the authors of this paper and was further optimized and adapted for efficient handling of verification queries within this work. Furthermore, since the decision procedure implemented in Astral is based on a translation of SL into SMT, Seal can seamlessly integrate SL reasoning with constraints from other theories.
As Seal is a relatively new tool, its potential has not yet been fully exploited. Currently, our analyser supports only three built-in data structures and simple integer properties. In the rest of the paper, we describe the architecture of the tool and its current capabilities. More details can be found in [7] (where a slightly earlier version of Seal is referred to as KTSN).
Abstract domain.
Seal represents abstract states using so-called symbolic heaps, i.e., SL formulae of the form $\exists x_{1},\ldots,x_{m}.\;\varphi_{1}*\ldots*\varphi_{n}$ where * denotes a separating conjunction and each $\varphi_{i}$ is an atomic formula. Atomic formulae are either pure describing (dis)equalities of pointers or constraints over non-pointer values, points-to atoms describing a single heap cell, or inductive predicates used to describe unbounded sequences of pointers of a particular shape. A special predicate is used to explicitly represent memory that was freed and is still reachable. The abstract domain can also represent precise values of integer variables within the range [k,k][-k,k] where $k=5$ by default. For each program location, a disjunction of symbolic heaps representing the reachable memory configurations is computed and represented as a set of symbolic heaps.
Currently, Seal supports three kinds of data structures: singly-linked lists (SLLs), doubly-linked lists (DLLs), and nested singly-linked lists (NLLs). In all three cases, the structure is acyclic and possibly empty. Some cyclic structures can still be represented using an explicit representation of a cyclic structure with multiple acyclic ones.
Abstraction.
Seal performs widening based on abstraction to generalise the shape of the memory by replacing its parts by instances of inductive predicates. The abstraction is performed by finding pairs of compatible spatial predicates (points-to or list predicates) and joining them into a single list predicate—provided that the memory location linking them is not the target of some external pointer links and it is not pointed to by a program variable. This way of abstraction is commonly used in other analysers, such as PredatorHP [13] or CPAchecker, too. The minimum length of a list predicate is tracked up to a pre-defined, fixed limit. For example, the formula $\exists y.\;\mathsf{ls}_{1^{+}}(x,y)*y\mapsto z*\psi$ (where $\mathsf{ls}_{n^{+}}(x,y)$ denotes a list segment from $x$ to $y$ of length greater than or equal to $n$) is abstracted to $\mathsf{ls}_{2^{+}}(x,z)*\psi$ provided that (1) $y$ does not occur in $\psi$ and (2) the original formula implies that $x\neq z$ (usually because $z$ is allocated in $\psi$) to guarantee acyclicity.
Currently, Seal guesses an expected inductive predicate for each C structure before the analysis. A structure with one or two self-referential fields is classified as an SLL or DLL, respectively. A structure with a single self-referential field and an additional field previously classified as an SLL is classified as an NLL. During the analysis, Seal just checks whether such an abstraction can be performed.
The above approach can fail in cases like the following structure TSLL defined as {struct TSLL* next; struct TSLL* inner;} and intended to be used to build an NLL with the next pointer unused in the nested lists. For this structure, Seal will guess its type to be DLL because it has two self-referential fields. This weakness could be improved by selecting the suitable inductive predicate for abstraction during the analysis. We would like to improve this aspect in the future together with adding more predicates for more data structures. We would also like to experiment with methods to synthesize inductive predicates from examples.
Finding loop fixpoints.
The analysis reaches a fixpoint for a loop header location when the entailment between the symbolic state $S$ from the current loop iteration and the state $T$ from the previous loop iteration is valid. The entailment is approximated by an incomplete method checking whether for each symbolic heap $S_{i}\in S$ there exists $T_{j}\in T$ such that $S_{i}\models T_{j}$. The pairs $(S_{i},T_{j})$ to be checked are chosen according to a heuristic based on the numbers of allocated memory locations in $S_{i}$ and $T_{j}$. For an entailment to be valid, there cannot be more allocations on the right side of the entailment than there are on the left side. The same method is used for pruning redundant symbolic heaps in the join operator. Additionally, a number of syntactic simplifications are applied to the states during analysis, such as reduction of equivalence classes in formulae, removal of variables outside of their scope, syntactic deduplication of formulae, etc.
Integer variables.
Currently, Seal can only track concretely known values of integer variables that are within a pre-configured range. Conditions and arithmetic operations on these variables can be evaluated precisely. When the result of an operation on such values goes out of the represented range, the variable that should receive the result becomes unconstrained instead. The range of tracked integer values can be configured, increasing the representable range trades precision for length of analysis.
Interprocedural analysis.
Analysis of function calls is optimised by storing function summaries – pairs of formulae representing precondition and postcondition of the analysed function. Using this information, a consecutive call of a function may be replaced by applying a matching summary.
Analysis results.
Seal is able to verify absence of NULL-pointer dereferences and use-after-free errors (valid-deref property), double-free errors (valid-free property), and some memory leaks (those that fall under the valid-memtrack property). The valid-memcleanup property is not yet supported. When an issue is found after reaching a condition that could not be precisely evaluated based on the current state (e.g., because of missing integer information), the analyser returns the UNKNOWN result. However, false positives are still possible because of the abstraction. In the future, they could be avoided by validating the issue by rerunning the analyser without abstraction. When an unsupported language feature is reached, the ERROR result is returned.
SL solver.
We briefly describe how the backend solver Astral of Seal handles SL queries. Given an SL formula, Astral first computes a bound on how many times it suffices to unfold the predicate instances present in the given SL formula and then uses this bound to translate the SL formula into several formulae in SMT theories of bitvectors and arrays. The translated formulae are then passed to an SMT solver, which is configured to be Bitwuzla [17] for SV-COMP. Unlike in [11], Astral runs in a mode supporting user-defined inductive predicates, i.e., the predicates are defined by Seal and Astral runs without any built-in support for them [10].
2 Strengths and Weaknesses
Table 1 provides a view on the results of MemSafety-LinkedLists [4], the only base category in which Seal participates, illustrating the strengths of Seal. Since almost half of the programs in this category manipulate lists of bounded size only and can therefore be verified using bounded model checking (or similar heuristics), we focus on the subset of benchmarks that require reasoning about unbounded data structures. Specifically, we exclude the benchmark sets list-simple and list-ext3-properties, which operate exclusively on bounded lists, as well as ddv-machzwd where the list-manipulating code is unreachable. Only the tools that can verify correctness of at least one of the remaining programs are shown in the table [^1] Since PredatorHP spuriously returns UNKNOWN for all tasks in SV-COMP’26, we used its results from SV-COMP’25. The benchmarks used for comparison were the same in both years..
Strengths.
As Table 1 shows, out of 23 verifiers participating in MemSafety, only 4 of them were able to verify correctness of programs with unbounded data structures. This indicates that reasoning about such programs remains a challenging task. Seal can verify more programs than 2ls [15, 14], but less than PredatorHP [13, 18] and CPAchecker [2, 1], two analysers based on symbolic memory graphs (SMG). We believe that the biggest strength of Seal lies in its potential to be extended for data structures that cannot be represented by SMGs, such as skip-lists or trees. Indeed, Astral in its new versions supports user-defined inductive predicates, including various kinds of trees, which may be implemented in Seal or potentially learned automatically. We would also like to use the generality of our technique to create a witness validator for correctness memory safety witnesses, once a witness format is established.
Weaknesses.
As an early prototype, Seal comes with a handful of limitations. First of all, we support only a subset of the C language (most kinds of pointer arithmetic, integer fields in list structures, global and static variables, arrays, and other language features are not supported). For this reason, we do not participate in other base categories of MemSafety. Seal currently has limited support for non-pointer values, only tracking precise values of integer variables, where known. For this reason, we are often not able to report a memory safety violation because an unknown condition in the program was reached, making any violation result a possible false positive. This limitation can be eliminated by tracking more complex information in symbolic states or by embedding our analyser as a new abstract domain into the Frama-C’s abstract interpreter EVA [5].
Table 1: The results of selected tools for MemSafety-LinkedLists and its subset with unbounded lists. The columns give numbers of correctly verified programs and found bugs, numbers of programs with no result reported (error, unknown result or out of resources), and the total time used for programs for which a result was produced.
All programs (134) Unbounded (69)
Analyser True False No res. Time [s] True False No res. Time [s]
PredatorHP 96 28 10 109 46 14 9 75
CPAchecker 87 27 20 2158 40 13 16 1550
SEAL 55 4 75 571 14 2 53 166
2LS 46 14 73 1142 3 13 53 96
3 Software Project and Tool Setup
Seal is implemented in the OCaml language as a plugin of the Frama-C platform [9, 3] (version 31.0). Currently, it is a stand-alone plugin that does not communicate with other analysers of Frama-C. Seal relies on the Frama-C’s intermediate representation (a fork of the CIL project [16]) and data flow analysis engine. Optionally, Seal can visualise its detailed results using the Frama-C’s new GUI frontend Ivette [8]. We used the Astral solver version 2.0-prerelease.1 that internally uses the Bitwuzla SMT solver version 0.8.2 through its OCaml binding.
Usage.
A binary of Seal is available at Zenodo [6] (the only runtime dependency is GCC). In SV-COMP, Seal is run via a Python3 script seal-entrypoint.py:
./seal-entrypoint.py -machdep=[gcc_x86_32|gcc_x86_64] input.c
where the machine models gcc_x86_32 and gcc_x86_64 correspond to the architectures ILP32 and LP64 used by SV-COMP, respectively. See the attached file README.md for more information.
Software project.
Seal is available under the MIT license, the project is maintained by Tomáš Brablec. The Astral solver which was modified as a part of this work is maintained by Tomáš Dacík.
Participation.
Seal participated in the MemSafety-LinkedLists base category only.
Data-Availability Statement.
Seal is available under the MIT license at https://github.com/pepega007xd/seal. The version participating in SV-COMP’26 is available under the tag svcomp26 and archived at Zenodo [6].
Acknowledgements.
We appreciate discussion with members of the Frama-C team at CEA, France. The research was supported by the Czech Science Foundation project 26-22640S. Tomáš Dacík was supported by the Brno PhD Talent Scholarship of the Brno City Municipality. The collaboration with CEA was supported by the project VASSAL: “Verification and Analysis for Safety and Security of Applications in Life” funded by the European Union under Horizon Europe WIDERA Coordination and support Action/Grant Agreement No. 101160022. [Uncaptioned image]